OpenStudy (winsteria):

PLEASE HELP IF YOU DO I'LL GIVE YOU MEDAL AND FAN please help!!! What is the value of the expression –c + d when c = 12.7 and d = –4.4? A. –8.6 B. –17.1 C. 8.6 D. 17.1

OpenStudy (nicoleced):

b

OpenStudy (kl0723):

-C+(-D)= -12.7+(-4.4)=-17.1 Therefore your answer is B :)

OpenStudy (winsteria):

Thank you both so much

OpenStudy (winsteria):

it's not letting me fan you guys idk why

OpenStudy (anonymous):

B